HOW MANY ML OF .1000 M SODIUM HYDROXIDE ARE REQUIRED TO TITRATE 25 ML OF .1502M OXALIC ACID
Answer #1

Moles of Oxalic acid = molarity×volume/1000

= 0.1502×25/1000 = 0.003755 moles

Moles of NaOH required = 2×moles of Oxalic acid

= 2×0.003755 = 0.00751 moles

Volume of NaOH = moles×1000/concentration

= 0.00751×1000/0.1000 ml = 75.1 ml

Please give feedback!